# RUIE: Retrieval-based Unified Information Extraction using Large Language Model

### v1

@@ Abstract
Unified information extraction (UIE) aims to complete all information extraction tasks using a single model or framework. While previous work has primarily focused on instruction-tuning large language models (LLMs) with constructed datasets, these methods require significant computational resources and struggle to generalize to unseen tasks. To address these limitations, we propose RUIE (Retrieval-based Unified Information Extraction), a framework that leverages in-context learning to enable rapid generalization while reducing computational costs. The key challenge in RUIE is selecting the most beneficial demonstrations for LLMs to effectively handle diverse IE tasks. To achieve this, we integrate LLM preferences for ranking candidate demonstrations and design a keyword-enhanced reward model to capture fine-grained relationships between queries and demonstrations. We then train a bi-encoder retriever for UIE through contrastive learning and knowledge distillation. To the best of our knowledge, RUIE is the first trainable retrieval framework for UIE. Experimental results on 8 held-out datasets demonstrate RUIE’s effectiveness in generalizing to unseen tasks, with average F1-score improvements of 19.22 and 3.13 compared to instruction-tuning methods and other retrievers, respectively. Further analysis confirms RUIE’s adaptability to LLMs of varying sizes and the importance of its key components.
@@
Xincheng Liao, Junwen Duanthanks:   Corresponding authors., Yixi Huang, Jianxin Wang Hunan Provincial Key Lab on Bioinformatics, School of Computer Science and Engineering, Central South University, Changsha, Hunan, China {ostars, jwduan, yx.huang}@csu.edu.cn, jxwang@mail.csu.edu.cn
@@ 1 Introduction
Figure 1: Illustration of 3 different paradigms for solving unified information extraction task.
@@
Unified Information Extraction (UIE) aims to employ one model or framework to extract diverse structured information from unstructured text, such as Named Entity Recognition (NER), Relationship Extraction (RE), and Event Extraction (EE). Compared to traditional task-specific information extraction (IE) methods, UIE exhibits superior generalization capabilities and higher practical utility Lu et al. (2022); Wang et al. (2023b); Li et al. (2024). Recently, large language models (LLMs) have shown impressive generalization abilities, achieving astonishing results across numerous tasks Wei et al. (2022); Zhang et al. (2024); Wang et al. (2024b). As a result, researchers are leveraging LLMs to tackle the challenges of UIE, which can be categorized into instruction-tuning based method and in-context learning based method (Figure 1).
@@
Due to the significant differences between the structured output of information extraction and large language model pre-training format, current large models still lag behind previous state-of-the-art approaches in IE Han et al. (2023); Ma et al. (2023); Chen et al. (2024). To bridge this gap, some researchers Wang et al. (2023b); Xiao et al. (2024b); Gui et al. (2024); Sainz et al. (2024); Li et al. (2024) have transformed annotated information extraction data into textual or code input-output pairs, training LLMs with instruction-tuning and achieving notable performance. However, instruction-tuning based method faces three main problems: (1) the high training cost, (2) potential loss of general capabilities Xu et al. (2024), (3) poor generalization and performance in unseen tasks.
In addition to instruction-tuning, large language models can also solve problems with just a few examples, which is known as in-context learning Brown et al. (2020). Li et al. (2023); Wang et al. (2023c) have shown that representing structured information in code format, combined with context learning, can outperform small fine-tuned UIE models in few-shot scenarios. Furthermore, Guo et al. (2023) enhances LLM extraction performance by applying offline sentence embedding to encode corpus and searching for relevant examples based on different input tasks. Moreover, some researchers have explored how to obtain more suitable semantic representations for information extraction tasks. Wang et al. (2023a); Wan et al. (2023) use fine-tuned smaller pre-trained language models to encode entities and relations as a retrieval index, achieving performance comparable to previous state-of-the-art methods. Nonetheless, current retrieval-based information extraction methods are mainly tailored to specific IE tasks, lacking UIE generalization ability, and only rely on semantic relevance for examples selection, ignoring the inherent preferences of LLMs.
To make full use of the generalization ability of LLMs and reduce computational costs, we introduce RUIE, a retrieval-based unified information extraction framework. RUIE aims to harness the model’s information extraction capabilities through in-context learning. As shown in Figure 1, compared to instruction tuning-based approaches, RUIE only requires fine-tuning a smaller dense retriever (million-level parameters), greatly reducing the fine-tuning cost. Furthermore, RUIE is compatible with various LLMs, acting as a versatile plugin. Compared to existing retrieval-based information extraction methods, we focus on “true” unified information extraction, where the candidate pool is a mixture of various information extraction tasks such as NER, RE and EE. RUIE is trained to select the most appropriate examples from candidate pool to optimize the LLM’s extraction performance. Furthermore, previous works solely rely on semantic relevance for example selection, we introduce LLM preferences to rank candidate demonstrations. In addition, considering the fine-grained nature of information extraction tasks Wang et al. (2023a); Wan et al. (2023), we have developed a keyword-enhanced reward model to focus on fine-grained information snippets beyond sentence-level semantics. In summary, our main contributions include:
  • We introduce RUIE, the first trainable retrieval framework for Unified Information Extraction (UIE), leveraging in-context learning to reduce computational costs while enabling rapid generalization to unseen tasks.
  • We propose a novel demonstration selection mechanism that incorporates LLM preferences for ranking candidates and utilizes a keyword-enhanced reward model to capture fine-grained relationships between queries and demonstrations.
  • RUIE demonstrates exceptional adaptability, effectively handling diverse IE tasks (NER, RE, EE) and showing compatibility with LLMs of varying sizes, thus serving as a universal plugin for UIE tasks.
@@ 2.1 Unified Information Extraction
Lu et al. (2022) first proposed the framework of unified Information Extraction. They used structured extraction language to unify the input and output forms of IE tasks and the structural schema instructor to guide the model generation, but it requires further fine-tuning for different downstream task. Lou et al. (2023) considers the knowledge transfer between different tasks and schemas, modeling UIE as a semantic matching task with three dimensions of token-token, token-label and lable-token, which achieves better zero/few-shot transfer ability. However, performing three-dimensional semantic matching for each token greatly increases the training and inference cost. Recently, some researchers have modeled IE as natural language Wang et al. (2023b); Gui et al. (2024) or code Sainz et al. (2024); Li et al. (2024) generation tasks. They constructed instruction-tuning datasets based on existing IE datasets to fine-tune large language models, realizing knowledge sharing of different tasks and effectively improving UIE performance. However, the fine-tuning cost of large language models is high, and the fine-tuned models do not generalize well in new domains.
@@ 2.2 In-context Learning based Information Extraction
In-context learning (ICL) is the ability of large language models to perform new tasks with only a few examples or demonstrations. One significant merit of ICL is the circumvention of fine-tuning, which might not always be possible due to limited access to the model parameters or constraints on computational resources Brown et al. (2020). Li et al. (2023); Wang et al. (2023c) represent structured IE tasks with codes, and improve extraction performance by randomly selecting examples. Guo et al. (2023) optimizes the random sample retrieval process by introducing sentence embedding to select sample according to the similarity between query and examples. Wang et al. (2023a); Wan et al. (2023) show that fine-grained alignment information such as entities and relations is more important than sentence similarity for example selection in IE tasks. They use the entity and relation representations obtained by the fine-tuned small model to replace the sentence representation for retrieval, and obtain better performance than sentence-level embedding. However, they must obtain the entity span in the sentence and the fine-tuned small model in advance, and the use scenario is limited. To the best of our knowledge, there is no trainable retrieval based framework designed for UIE. On the one hand, we do not need any prior information about the sentence, on the other hand, we design a retrieval training scheme for UIE, which achieves better performance than general sentence embedding.
@@ 3.1 Problem Statement
Figure 2: The overall architecture of RUIE. The training process consists of three steps: 1) the sparse retriever bm25 initializes a candidate set, which is then scored by the LLM. 2) a keyword-enhanced reward model captures fine-grained information. 3) a bi-encoder dense retriever is trained using contrastive learning and knowledge distillation. During inference, the trained dense retriever selects the best demonstrations from candidate pool $P$, and passes them to the LLM to produce the output.
@@ 3.2 LLM Preference Scoring
Sample Format. In the candidate pool $P$, each task sample $(x_{i},y_{i})$ consists of four parts: (1) Task Name: name of a specific IE task, such as “Named Entity Recognition”. (2) Schema: task ontology presented in the form of python list. (3) Input: input context to be extracted. (4) Output: structured output linearized by natural language, such as “Entitytype1: EntityName1; …”.
Score examples using LLM. Previous retrieval based information extraction methods usually only consider the text similarity between the input text and the candidate pool examples, ignoring the preference of the LLM. We assume that the LLM knows what good examples are, specifically, good examples can maximize the probability that the model outputs ground-truth $y_{truth}$. Therefore, given an input $s=(x,y)$ in the training set, we enumerate each example $e_{i}=(x_{i},y_{i})$ in the candidate pool $P$. To keep with the inference stage, we concatenate the extraction instruction $I$ (see Appendix D), example $e_{i}$, and input $s$ into the model. We use the token-level average log-likelihood of $y$ as the score between input $s$ and example $e_{i}$:
$Score(s,e_{i})=log~{}p(y|I;x_{i};y_{i};x)$ (1)
Then we rank the all candidate examples in descending order and select top $k$ and last $n$ examples as positive and negative examples respectively.
Initialize candidates. In practice, the number of examples in the candidate pool $P$ is usually very large. To reduce the costs of LLM scoring, given an input sample $s=(x,y)$, we use a sparse retriever bm25 Robertson and Zaragoza (2009) to retrieve top-k candidate samples $\{(x_{i},y_{i})\}_{i}^{k}$ in candidate pool $P$ as the set of scoring candidates. Choosing an appropriate k value can greatly improve the scoring efficiency and ensure that the candidate set contains both positive and negative samples.
@@ 3.3 Keyword-enhanced Reward
Some studies have shown that the alignment of query and candidate examples on fine-grained information such as entities and relations is more important than the alignment on coarse-grained information, like sentence semantics. Therefore, in order to make the fine-grained information between the input query and the candidate examples fully interactive, we propose a keyword-enhanced training strategy based on cross-encoder. On the one hand, the strategy aligns the fine-grained information through keyword, and on the other hand, the cross-encoder realizes the full interaction between query and candidate examples by accessing the ground-truth. Specifically, given a training sample $(x,y)$, for each information snippets $(s,o)$ in $y$, where $o$ is the label of span $s$, we add a pair of special tags “<Keyword>” and “</Keyword>” around $s$ in context $x$ (Figure 2). Given an enhanced input $(x^{\prime},y^{\prime})$, we sample a positive example $(x^{\prime}_{+},y^{\prime}_{+})$ in the top-k of the enhanced ranked candidates and take the last-n as the negative examples $(x^{\prime}_{-_{i}},y^{\prime}_{-_{i}})$. We train the cross-encoder using the cross-entropy loss:
@@ 3.4 UIE Retriever Training
In the test phase, we have no access to the ground-truth output $y$ corresponding to the input $x$. So we cannot directly use the above cross-encoder as the retriever. To improve retrieval efficiency, we build a UIE retriever based on a bi-encoder. Specifically, we encode all samples $e^{\prime}_{i}=(x^{\prime}_{i},y^{\prime}_{i})$ in the keyword-enhanced candidate pool $P$ into vector $h_{e^{\prime}i}$ and build the index in advance. Given an input $x$, we use the same encoder to encode it into a vector $h_{x}$ and calculate matching score between the input $x$ and each example $e^{\prime}_{i}$ using cosine similarity:
$Score(x,e^{\prime}_{i})=\cos{(h_{x},h_{e^{\prime}i})}$ (3)
$h_{e^{\prime}i}$ $=\text{avgpool}(Encoder(x^{\prime}_{i};y^{\prime}_{i}))$
@@
where “avgpool” refers to average pooling and “;” refers to concatenation operation.
We use two kinds of supervision signals to train the UIE retriever: (1) In order to make full use of the positive and negative pairs discovered by LLM, we use the Info-NCE loss $L_{contrastive}$ to perform contrastive learning between the positives and in-batch negatives. (2) In order to make full use of the fine-grained alignment information of the reward model, we use the KL divergence to align the output distributions of the reward model and the retriever, calculated as $L_{distill}$, and the final training loss is the weighted sum of the above two losses:
@@
where $\alpha$ is a hyper-parameter to measure the importance of the above two losses.
@@
Table 1: Performance (in F1%) on NER, RE, ED and EAE tasks under held-out settings. Bold indicates the highest scores and the second-best scores are underlined. The only difference between RUIE and other retrieval-based methods is that they use different retrievers. It’s worth noting that the model size of InstuctUIE and LLaMA2-IEPILE are 11B and 13B, respectively, while the default model size of RUIE is 8B.
@@ 4.1 Datasets
To exhaustively evaluate the generalization ability of RUIE, we collect 31 held-in and 8 held-out datasets. We used the training set of the held-in and held-out datasets to form the candidate pool $P$. We constructed the retriever training set based on the held-in dataset, specifically, we sampled 10000 samples from each dataset and included all examples from datasets with less than 10000 samples. The Held-out dataset samples was completely unseen during training, and the UIE Retriever was tested on the test set of held-in and held-out data after training.

### v2

@@ Abstract
Unified information extraction (UIE) aims to extract diverse structured information from unstructured text. While large language models (LLMs) have shown promise for UIE, they require significant computational resources and often struggle to generalize to unseen tasks. We propose RUIE (Retrieval-based Unified Information Extraction), a framework that leverages in-context learning for efficient task generalization. RUIE introduces a novel demonstration selection mechanism combining LLM preferences with a keyword-enhanced reward model, and employs a bi-encoder retriever trained through contrastive learning and knowledge distillation. As the first trainable retrieval framework for UIE, RUIE serves as a universal plugin for various LLMs. Experimental results on eight held-out datasets demonstrate RUIE’s effectiveness, with average F1-score improvements of 19.22 and 3.22 compared to instruction-tuning methods and other retrievers, respectively.
@@
Xincheng Liao, Junwen Duanthanks:   Corresponding author. Email: jwduan@csu.edu.cn, Yixi Huang, Jianxin Wang Hunan Provincial Key Lab on Bioinformatics, School of Computer Science and Engineering, Central South University, Changsha, Hunan, China {ostars, jwduan, yx.huang}@csu.edu.cn, jxwang@mail.csu.edu.cn https://github.com/OStars/RUIE
@@ 1 Introduction
Figure 1: Illustration of three different paradigms for solving unified information extraction task.
@@
Unified Information Extraction (UIE) represents a paradigm shift from traditional task-specific approaches, aiming to extract diverse structured information (e.g., Named Entity Recognition, Relationship Extraction, and Event Extraction) using a single model or framework. This unified approach demonstrates superior generalization capabilities and practical utility compared to conventional methods that require separate models for different extraction tasks Lu et al. (2022); Wang et al. (2023b); Li et al. (2024). With the emergence of large language models (LLM) and their remarkable generalization abilities in various tasks Wei et al. (2022); Zhang et al. (2024); Wang et al. (2024b); Jia et al. (2024), researchers have begun to explore LLM-based solutions to UIE challenges through two main approaches: instruction tuning and in-context learning (Figure 1).
@@
However, the inherent mismatch between structured information extraction outputs and LLM’s pretraining format poses significant challenges, resulting in current large models still underperforming specialized IE approaches Han et al. (2023); Ma et al. (2023); Chen et al. (2024). Although some researchers Wang et al. (2023b); Xiao et al. (2024b); Gui et al. (2024); Sainz et al. (2024); Li et al. (2024) have bridged this gap by transforming IE annotations into textual or code pairs for instruction tuning, this approach faces several critical limitations: substantial computational costs, potential degradation of general capabilities Xu et al. (2024), and limited generalization to unseen tasks.
Alternatively, in-context learning Brown et al. (2020) has emerged as a promising direction, allowing LLMs to perform tasks with minimal examples or demonstrations. Recent works have demonstrated its effectiveness in IE tasks through various approaches: representing structured information in code format Li et al. (2023); Wang et al. (2023c), improving extraction through offline sentence embedding and example retrieval Guo et al. (2023), and developing task-specific semantic representations Wang et al. (2023a); Wan et al. (2023). Despite these advances, existing retrieval-based methods remain largely task-specific, lacking true UIE capabilities, and primarily relying on semantic relevance while overlooking LLMs’ inherent preferences in example selection.
To address these limitations while leveraging LLMs’ capabilities, we propose RUIE, a novel retrieval-based unified information extraction framework. As illustrated in Figure 1, RUIE significantly reduces computational costs compared to instruction-tuning approaches by only requiring the fine-tuning of a smaller dense retriever (million-level parameters). Unlike existing retrieval-based methods, RUIE achieves true unified information extraction by maintaining a diverse candidate pool spanning multiple IE tasks (NER, RE, EE) and incorporating both semantic relevance and LLM preferences in example selection. Furthermore, we introduce a keyword-enhanced reward model to capture the label and fine-grained information, addressing the detailed nature of IE tasks Wang et al. (2023a); Wan et al. (2023); Duan et al. (2024a, b). Our main contributions are as follows.
  • We propose RUIE, a trainable retrieval framework for UIE that enables efficient task generalization through in-context learning while reducing computational costs.
  • We develop an innovative demonstration selection mechanism that uniquely combines LLM preferences with a keyword-enhanced reward model, enabling more accurate and context-aware example selection.
  • We demonstrate RUIE’s strong flexibility as a general IE framework, showing robust performance across different tasks (NER, RE, EE) and easy integration with various LLM architectures.
@@ 2.1 Unified Information Extraction
Lu et al. (2022) first proposed the framework of unified Information Extraction. They used structured extraction language to unify the input and output forms of IE tasks and the structural schema instructor to guide the model generation, but it requires further fine-tuning for different downstream tasks. Lou et al. (2023) considers the knowledge transfer between different tasks and schemas, modeling UIE as a semantic matching task with three dimensions of token-token, token-label and label-token, which achieves better zero/few-shot transfer ability. However, performing three-dimensional semantic matching for each token greatly increases the training and inference cost. Recently, some researchers have modeled IE as natural language Wang et al. (2023b); Gui et al. (2024) or code Sainz et al. (2024); Li et al. (2024) generation tasks. They constructed instruction-tuning datasets based on existing IE datasets to fine-tune large language models, realizing knowledge sharing of different tasks and effectively improving UIE performance. However, fine-tuning large language models is expensive, and the fine-tuned models do not generalize well in new domains.
Figure 2: The overall architecture of RUIE. The training process consists of three steps: 1) the sparse retriever bm25 initializes a candidate set, which is then scored by the LLM. 2) a keyword-enhanced reward model captures fine-grained information. Keyword-enhanced strategy only applies to the input field of the example. 3) a bi-encoder dense retriever is trained using contrastive learning and knowledge distillation. During inference, the trained dense retriever selects the best demonstrations from the candidate pool $P$, and passes them to the LLM to produce the output.
@@ 2.2 In-context Learning based Information Extraction
In-context learning (ICL) is the ability of large language models to perform new tasks with only a few examples or demonstrations. One significant merit of ICL is the circumvention of fine-tuning, which might not always be possible due to limited access to the model parameters or constraints on computational resources Brown et al. (2020). Li et al. (2023); Wang et al. (2023c) represent structured IE tasks with codes, and improve extraction performance by randomly selecting demonstrations. Guo et al. (2023) optimizes the random retrieval process by introducing sentence embedding to select demonstrations according to the similarity between query and demonstrations. Wang et al. (2023a); Wan et al. (2023) show that fine-grained alignment information such as entities and relations is more important than sentence similarity for example selection in IE tasks. They use the entity and relation representations obtained by the fine-tuned small model to replace the sentence representation for retrieval, and obtain better performance than sentence-level embedding. However, they must obtain the entity span in the sentence and the fine-tuned small model in advance, which restricts the use cases. To the best of our knowledge, there is no trainable retrieval-based framework designed for UIE. On the one hand, we do not need any prior information about the input sentence, on the other hand, we design a retrieval training scheme for UIE, which achieves better performance than general sentence embedding.
@@ 3.2 LLM Preference Scoring
Sample Format. In the candidate pool $P$, each task sample $(x_{i},y_{i})$ consists of four parts: (1) Task Name: name of a specific IE task, such as “Named Entity Recognition”. (2) Schema: task ontology presented in the form of python list. (3) Input: input context to be extracted. (4) Output: structured output linearized by natural language, such as “Entitytype1: EntityName1; …”. (see Appendix D)
Score examples using LLM. Previous retrieval-based information extraction methods usually only consider the text similarity between the input text and the candidate pool examples, ignoring the preference of the LLM. We assume that the LLM knows what good examples are; specifically, good examples can maximize the probability that the model produces ground-truth $y_{truth}$. Therefore, given an input $s=(x,y)$ in the training set, we enumerate each example $e_{i}=(x_{i},y_{i})$ in the candidate pool $P$. To keep with the inference stage, we concatenate the extraction instruction $I$ (see Appendix E), example $e_{i}$, and input $s$ into the model. We serve the token-level average log-likelihood of $y$ as the score between input $s$ and example $e_{i}$:
$Score(s,e_{i})=\log~{}p(y|I;x_{i};y_{i};x)$ (1)
Then we rank the all candidate examples in descending order and select top $k$ and last $n$ as positive and negative examples respectively.
Initialize candidates. In practice, the candidate pool $P$ often comprises a significantly large number of examples. To reduce the costs of LLM scoring, given an input sample $s=(x,y)$, we use a sparse retriever bm25 Robertson and Zaragoza (2009) to retrieve top-k candidate samples $\{(x_{i},y_{i})\}_{i}^{k}$ in the candidate pool $P$ as the set of scoring candidates. Choosing an appropriate k can greatly improve the scoring efficiency and ensure that the candidate set contains both positive and negative samples.
@@ 3.3 Keyword-enhanced Reward
Wang et al. (2023a); Wan et al. (2023) have shown that the alignment of query and candidate examples on fine-grained information such as entities and relations is more important than the alignment on coarse-grained information, like sentence semantics. Therefore, in order to make the fine-grained information between the input query and the candidate examples fully interactive, we propose a keyword-enhanced training strategy based on cross-encoder. On the one hand, the strategy aligns the fine-grained information through keyword, and on the other hand, the cross-encoder realizes the full interaction between query and candidate examples by accessing the ground-truth. Specifically, given a training sample $(x,y)$, for each information snippet $(sp,o)$ in $y$, where $o$ is the label of span $sp$, we add a pair of special tags “<Keyword>” and “</Keyword>” around $sp$ in context $x$ (Figure 2). Given an enhanced input $(x^{\prime},y^{\prime})$, we sample a positive example $(x^{\prime}_{+},y^{\prime}_{+})$ in the top-k of the enhanced ranked candidates, take the last-n as the negative examples $(x^{\prime}_{-_{i}},y^{\prime}_{-_{i}})$ and produce a real-valued score $s$. We train the cross-encoder using the cross-entropy loss:
@@ 3.4 UIE Retriever Training
In the test phase, we have no access to the ground-truth output $y$ corresponding to the input $x$. So we cannot directly use the above cross-encoder as the retriever. To enhance retrieval efficiency, we construct UIE retriever based on bi-encoder architecture. Specifically, we encode all samples $e^{\prime}_{i}=(x^{\prime}_{i},y^{\prime}_{i})$ in the keyword-enhanced candidate pool $P$ into vector $h_{e^{\prime}i}$ and build the index in advance. Given an input $x$, we use the same encoder to encode it into a vector $h_{x}$ and calculate matching score between the input $x$ and each example $e^{\prime}_{i}$ using temperature-scaled dot product:
$score(x,e^{\prime}_{i})=h_{x}\cdot h_{e^{\prime}i}/\tau$ (3)
$h_{e^{\prime}i}$ $=\text{avgpool}(Encoder(x^{\prime}_{i},y^{\prime}_{i}))$
@@
where “avgpool” refers to average pooling and $\tau$ refers to temperature hyperparameter.
We use two kinds of supervision signals to train the UIE retriever: (1) In order to make full use of the positive and negative pairs discovered by LLM, we use the Info-NCE loss $L_{contrastive}$ to perform contrastive learning between the top-k positives and in-batch negatives. (2) In order to make full use of the fine-grained alignment information of the reward model, we use the KL divergence to align the output distributions of the reward model and the retriever, calculated as $L_{distill}$, and the final training loss is the weighted sum of the above two losses:
@@
where $\alpha$ is a hyperparameter to measure the importance of the above two losses.
@@
Table 1: Performance (in F1-score) on NER, RE, ED and EAE tasks under held-out settings. Bold indicates the highest scores and the second-best scores are underlined. “BM25*” and “Random*” refer to using BM25 or randomly selecting demonstrations from the candidate pool corresponding to the task, rather than from the multi-task candidate pool. The only difference between RUIE and other retrieval-based methods is that they use different retrievers. It’s worth noting that the model size of InstuctUIE and LLaMA2-IEPILE are 11B and 13B, respectively, while the default model size of RUIE is 8B.
@@ 4.1 Datasets
To exhaustively evaluate the generalization ability of RUIE, we collect 31 held-in and 8 held-out datasets. We used the training set of the held-in and held-out datasets to form the candidate pool $P$. We constructed the retriever training set based on the held-in dataset, specifically, we sampled 10000 samples from each dataset and included all examples from datasets with less than 10000 samples. The Held-out dataset samples was completely unseen during training, and the UIE Retriever was tested on the test set of held-in and held-out datasets after training.